# On Topology of Three-dimensional Continua

### v2

@@ Abstract
We propose to model the topology of three-dimensional (3D) continua by Yin sets, regular open semianalytic sets with bounded boundary. Our model differs from manifold-based models in that singular points of a 3D continuum, i.e., boundary points where the tangent plane is not uniquely defined, are treated not as anomalies but as a central subject of our theoretical investigation. We characterize the local and global topology of Yin sets. Then we give a unique boundary representation of Yin sets based on the notion of a glued surface, a quotient space of an orientable compact 2-manifold along a one-dimensional CW complex. Our results apply to 3D continua with arbitrarily complex topology and may be useful in a number of scientific and engineering applications such as solid modeling, computer-aided design, and numerical simulations of multiphase flows with topological changes.
@@ 1.4 Contributions of this work
The above discussions sum up to questions as follows.
@@
    Based on different types of intersections in Definition 4.2 and the answer to (Q-2), can we uniquely represent each 3D continuum by a set of surfaces that are free of overlaps, self-intersections, and proper pairwise intersections?
In this work, we give positive answers to all the above questions by generalizing Yin sets from 2D to 3D; this is also our first step in 3D fluid modeling.
In Section 2, we briefly review the Yin space proposed in [34]; it is fortuitous that the Yin space originally designed for 2D continua worked out in 3D without any modifications. The 3D Yin space is also our answer to (Q-1). In Section 3, we define several types of singular points to include all characteristic boundary points of a 3D Yin set. In Section 4, we define a surface as an orientable compact 2-manifold and introduce the important notions of proper and improper intersections of surfaces. In Section 5, we examine the local topology of 3D continua by focusing on the good neighborhood of singular points. The key concept of good pairing in Section 5.4 leads to Theorem 6.3 on the global topology, which answers (Q-2). Different from the 2D case, it is not obvious how to deduce a unique B-rep of 3D continua from the main theorem on the global topology. In Section 7, we define the new concept of glued surfaces, employ it as the basic building block in decomposing the boundary of a 3D continuum, and show in Theorem 7.11 that any connected 3D Yin set can be uniquely represented by a set of oriented glued surfaces. Apart from answering (Q-3), this unique 3D B-rep has the same form as that of the 2D case, with glued surfaces replacing the Jordan curves in [34, eq. (3.6)]. Finally in Section 8, we conclude this paper with several future research prospects.
@@ 7 Boundary representation
Theorem 7.11 is applied to each connected component of a general Yin set to get

### v3

@@ Abstract
We propose to model the topology of three-dimensional (3D) continua by Yin sets, regular open semianalytic sets with bounded boundary. Our model differs from manifold-based models in that singular points of a 3D continuum, i.e., boundary points where the tangent plane is not uniquely defined, are treated not as anomalies but as a central subject of our theoretical investigation. We characterize the local and global topology of Yin sets. Then we give a unique boundary representation (B-rep) of Yin sets based on the notion of a glued surface, a quotient space of an orientable compact 2-manifold along a one-dimensional CW complex. As a prominent advantage of this B-rep, global topological invarants such as the number of connected components and the number of holes (2-cycles) in each component can be extracted in $O(1)$ time. Our results apply to 3D continua with arbitrarily complex topology and may be useful in a number of scientific, engineering, and industrial applications such as solid modeling, computer-aided design, and numerical simulations of multiphase flows with topological changes.
@@ 1.4 Contributions of this work
The discussions in Sections 1.1, 1.2, and 1.3 manifest common topological and geometric problems in a wide range of scientific, engineering, and industrial endeavors.
@@
    Based on the answer to (Q-2), can we design a scheme of unique B-rep of 3D continua by a set of surfaces that are free of overlaps, self-intersections, and proper pairwise intersections?
  4. (Q-4)
    Sometimes the physics of a 3D continuum depends on global topological invariants such as the number of connected components of a fluid phase and the number of holes inside a connected component. For example, the number of bubbles per unit volume in a bubbly flow is a crucial characterization of the fluid properties. So can we extract these topological invariants in $O(1)$ time?
In this work, we give positive answers to all the above questions.
In Section 2, we briefly review the Yin space proposed in [34]; it is fortuitous that the Yin sets in Definition 2.2 originally designed for 2D continua worked out in 3D without any modifications. The 3D Yin space is also our answer to (Q-1). In Section 3, all characteristic boundary points of a 3D Yin set are included in the notion of singular points. In Section 4, we define a surface as an orientable compact 2-manifold and introduce the important notions of proper and improper intersections of surfaces. In Section 5, we examine the local topology of 3D continua by focusing on the good neighborhood of singular points. The key concept of good pairing in Section 5.4 leads to Theorem 6.3 on the global topology, which answers (Q-2). Different from the 2D case, it is not obvious how to deduce a unique B-rep of 3D continua from the main theorem on the global topology. In Section 7, we define the new concept of glued surfaces, employ it as the basic building block in decomposing the boundary of a 3D continuum, and show in Theorem 7.11 that any connected 3D Yin set can be uniquely represented by a set of oriented glued surfaces. Apart from answering (Q-3,4), this unique 3D B-rep has the same form as that of the 2D case, with glued surfaces replacing the Jordan curves in [34, eq. (3.6)].
Following the theory [34] and algorithms [25] of 2D fluid modeling, this work is our first theoretical step in 3D fluid modeling for multiphase flows.
@@ 7 Boundary representation
Theorem 7.11 is applied to each connected component of a general Yin set to give
@@
Corollary 7.13 furnishes an appealing advantage of the B-rep that global topological invariants of any 3D Yin set $\cal Y$} can be extracted in $O(1)$ time. In particular, the number of connected components of $\cal Y$} is 1 if ${\cal G}_{\partial{\cal Y}}$ is of the type $\cal G$$^{-}$; otherwise it is the number of positively oriented glued surfaces in ${\cal G}_{\partial{\cal Y}}$. Also, the number of holes (i.e., 2-cycles) in the closure $\overline{{\cal Y}_{i}}$ of a connected component of $\cal Y$} is the number of negatively oriented glued surfaces in ${\cal G}_{\partial{\cal Y}_{i}}$.